OpenStudy (anonymous):

Find the slope and y-intercept of each line. y = x – 3 A. slope: 3, y-intercept: –1 B. slope: –3, y-intercept: 1 C. slope: 1, y-intercept: –3 D. slope: –1, y-intercept: –3

OpenStudy (anonymous):

So the slope is the number before x....

OpenStudy (anonymous):

so what number is that

OpenStudy (anonymous):

y

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

3

OpenStudy (anonymous):

i mean like: y=mx+b so what is m in this case?

OpenStudy (anonymous):

idk???

OpenStudy (anonymous):

ok here this problem might be a little confusing..... so take this for example: y=3x+5

OpenStudy (anonymous):

the number before x is 3

OpenStudy (anonymous):

that's is your slope

OpenStudy (anonymous):

y=mx+b m= slope

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

So in this case:y=x-3 so there's no number before x so it's one. because x times 1 is still x.....

OpenStudy (anonymous):

does that help a little more?

OpenStudy (anonymous):

so C is the answer???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because in the equation: y=mx+b b is the y intercept...

OpenStudy (anonymous):

did that help at all?

OpenStudy (anonymous):

im soo sorry u have to put up with this...........if u want u dont have to help me ill still give u best response

OpenStudy (anonymous):

no its totally fine

OpenStudy (anonymous):

is C the answer!!!!

OpenStudy (anonymous):

yes!!!!

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

ahhhhhhhhhh!!!!!!! OMG,ur a legend!!!! :)

OpenStudy (anonymous):

thanks! i got to go now.... hope that helped a lot :-) if you ever have a q just ask me.

OpenStudy (anonymous):

kk,i totally will:)

OpenStudy (anonymous):

bye